# ShoeModel: Learning to Wear on the User-specified Shoes via Diffusion Model

### v1

@@ 1 Introduction
With the help of the pre-trained stable diffusion model [17], many multi-modal conditional image synthesis methods [25, 12, 9] have ability to generate controllable images. These methods aim to precisely control the generation process. However, these methods cannot preserve the identity of the objects in the conditional images which is crucial for the synthesis of realistic electronic business advertising images. Moreover, the interaction between conditional objects and synthesis human is also not considered in these existing methods.
@@ 2 Related Works
In addition, we noticed that both training-based methods like DreamBooth [18] and training-free methods like IP-adapter [24] and Anydoor [3], claim to be capable of keeping the identity of specified objects. However, most of them mainly show the examples of person or animals or cartoon human which have one thing in common, i.e., these are visually insensitive to subtle visual changes. When being applied in e-commerce scenario, their performances are worse since the identity of the given object cannot be exactly guaranteed. In contrast, painting-based methods like Paint-by-example [23] and ControlNet-inpaint [25] will always generate something new contents around the given objects and cannot produce reasonable interactions with human. All the above are verified in section 5.2.
@@ 3.1 Wearable-area Detection Module
Figure 3: Visualization of the visible area and the wearable area. The visible area (shown in blue) can be seen in both (2(a)) and (2(b)), while the wearable area (shown in yellow) only can be seen in (2(a)).
@@ 3.3 Shoe-wearing Module
Shoe Image Area Annot. Shoe Image Area Annot.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 4.1 Wearable-area detection dataset
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 5 Experiments
FID $\downarrow$ KID×${}_{\times 1k}$downarrow CLIP score $\uparrow$ Specified-Shoe Quality of Generation Num $\uparrow$
@@ 5.2 Qualitative Analysis
Figure 5: Qualitative comparisons with SDXL [14], Paint-by-example [23], ControlNet-Inpainting [25], AnyDoor [3] and Ip-Adapter [24]. One can observe that our ShoeModel performs the best in terms of image quality, shoe-ID consistency and reasonability of interactions between human and shoes. More results are in supplementary file.
Since the quantitative results sometimes diverge lots with human preference, we provide the qualitative comparisons with several representative related works, including text2img-based SDXL [14], painting-based Paint-by-example [23] and ControlNet-Inpainting [25], training-free based AnyDoor [3] and Ip-Adapter [24], and training-based Dreambooth [18]. For fair comparisons, we use the same input shoes and prompts for all methods. From Figure 5, we observe that the T2I-based SDXL[14] naturally doesn’t support interactions with the given shoes, since it only receives text input.
@@
Compared with Training-free methods: different from [23, 25], AnyDoor [3] and Ip-Adapter [24] first compress the input objects into high-level feature embeddings via certain image encoders, and then produce output image conditioned on these image embeddings. However, these kind of methods cannot guarantee the ID consistency between input and output. It is because both of them use an image encoder (DINOv2 or CLIP) to compress the pixel-level shoes into high-level latent-space and then generate images based on this latent-space, the image encoder is pre-trained to mainly focus on the object-level concepts instead of the minor details within objects, such as pixel-level color, logo, texture and so on. As a result, information loss of shoe-detail is inevitable during this compression process, which in turn affects the consistency of the generated shoe-ID. Figure 5.(e,f) verifies this.
@@ 6 Discussion
Impact: Although this paper focuses on a specific task, i.e., how to wear on the given shoes when exactly maintaining its ID(including color, shape, logo and any other things from the given shoes). There is no existing method can conduct this task. And This application is very important in e-commerce system since "shoes" accounts for a large proportion in online shopping. If we can automatically generate realistic advertising images, the marketing costs will be significantly decreased since there is no longer a need to hire real-life models to wear the given shoes for shooting promotional product images of those shoes.
Moreover, our pipeline might inspire others for handling other similar tasks like: how to grasp the object with hands, how to interact with other products like hat, glasses and clothes when exactly maintaining the object identity. And our paper brings new ideas and perspectives to the field of strictly preserving IDs, instead of blindly following the training-based or training-free pipelines.

### v2

@@ 1 Introduction
With the help of the pre-trained stable diffusion model [17], many multi-modal conditional image synthesis methods [26, 12, 9, 27, 23, 3] have ability to generate controllable images. These methods aim to precisely control the generation process. However, these methods cannot preserve the identity of the objects in the conditional images which is crucial for the synthesis of realistic electronic business advertising images. Moreover, the interaction between conditional objects and synthesis human is also not considered in these existing methods.
@@ 2 Related Works
In addition, we noticed that both training-based methods like DreamBooth [18] and training-free methods like IP-adapter [25], claim to be capable of keeping the identity of specified objects. However, most of them mainly show the examples of person or animals or cartoon human which have one thing in common, i.e., these are visually insensitive to subtle visual changes. When being applied in e-commerce scenario, their performances are worse since the identity of the given object cannot be exactly guaranteed. In contrast, painting-based methods like Paint-by-example [24] and ControlNet-inpaint [26] will always generate something new contents around the given objects and cannot produce reasonable interactions with human. All the above are verified in section 5.2.
@@ 3.1 Wearable-area Detection Module
Figure 3: Visualization of the visible area and the wearable area. The visible area (shown in blue) can be seen in both (3(a)) and (3(b)), while the wearable area (shown in yellow) only can be seen in (3(a)).
@@ 4.1 Wearable-area detection dataset
Shoe Image Area Annot. Shoe Image Area Annot.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 5 Experiments
FID $\downarrow$ KID×1k $\downarrow$ CLIP score $\uparrow$ Specified-Shoe Quality of Generation Num $\uparrow$
@@ 5.2 Qualitative Analysis
Figure 5: Qualitative comparisons with SDXL [14], Paint-by-example [24], ControlNet-Inpainting [26] and Ip-Adapter [25]. One can observe that our ShoeModel performs the best in terms of image quality, shoe-ID consistency and reasonability of interactions between human and shoes. More results are in supplementary file.
Since the quantitative results sometimes diverge lots with human preference, we provide the qualitative comparisons with several representative related works, including text2img-based SDXL [14], painting-based Paint-by-example [24] and ControlNet-Inpainting [26], training-free based Ip-Adapter [25], and training-based Dreambooth [18]. For fair comparisons, we use the same input shoes and prompts for all methods. From Figure 5, we observe that the T2I-based SDXL[14] naturally doesn’t support interactions with the given shoes, since it only receives text input.
@@
Compared with Training-free methods: different from [24, 26], Ip-Adapter [25] first compresses the input objects into high-level feature embeddings via certain image encoders, and then produces output image conditioned on these image embeddings. However, this kind of method cannot guarantee the ID consistency between input and output. It is because it uses an image encoder (CLIP) to compress the pixel-level shoes into high-level latent-space and then generates images based on this latent-space, the image encoder is pre-trained to mainly focus on the object-level concepts instead of the minor details within objects, such as pixel-level color, logo, texture and so on. As a result, information loss of shoe-detail is inevitable during this compression process, which in turn affects the consistency of the generated shoe-ID. Figure 5.(e,f) verifies this.
@@
BG Visable-Area Wearable-Area All
IoU Acc IoU Acc IoU Acc mIoU mAcc aAcc
WD 99.42 99.67 97.3 98.62 85.55 92.45 94.09 96.91 98.60
Other-Area Wearable-Area All
IoU Acc IoU Acc mIoU mAcc aAcc
WD(2-classes) 99.45 99.60 79.59 89.15 89.52 94.38 98.11
@@ 6 Discussion
Acknowledgement
This work was supported in part by the National Key R&\&&D Program of China (2021YFF0900500), and the National Natural Science Foundation of China (NSFC) under grants 62441202.